# A Starting Point for Dynamic Community Detection with Leiden Algorithm

### v1

@@ A.1. Our Parallel Naive-dynamic (ND) Leiden
Algorithm 1 details our multicore implementation of Naive dynamic (ND) Leiden. Here, vertices are assigned to communities based on the previous snapshot of the graph, with all vertices being processed regardless of edge deletions and insertions in the batch update. The algorithm accepts the previous graph snapshot $G^{t-1}$, the current graph snapshot $G^{t}$, edge deletions $\Delta^{t-}$ and insertions $\Delta^{t+}$ from the batch update, the previous community membership $C^{t-1}$ for each vertex, the weighted degree of each vertex $K^{t-1}$, and the total edge weight of each community $\Sigma^{t-1}$. The output includes the updated community memberships $C^{t}$, the updated weighted degrees $K^{t}$, and the updated total edge weights of communities $\Sigma^{t}$.
@@
A.2. Our Parallel Delta-screening (DS) Leiden
Algorithm 2 presents the pseudocode for our multicore implementation of Delta-screening (DS) Leiden. It employs modularity-based scoring to identify an approximate region of the graph where vertices are likely to change their community membership (Zarayeneh and Kalyanaraman, 2021). It takes as input the previous graph snapshot $G^{t-1}$, the current snapshot $G^{t}$, edge deletions $\Delta^{t-}$ and insertions $\Delta^{t+}$ from the batch update, the previous community memberships of vertices $C^{t-1}$, the weighted degrees of vertices $K^{t-1}$, and the total edge weights of communities $\Sigma^{t-1}$. The algorithm outputs the updated community memberships $C^{t}$, weighted degrees $K^{t}$, and total edge weights of communities $\Sigma^{t}$. Prior to processing, the batch update — which includes edge deletions $(i,j,w)\in\Delta^{t-}$ and insertions $(i,j,w)\in\Delta^{t+}$ — is sorted separately by the source vertex ID $i$.
@@ A.2. Our Parallel Delta-screening (DS) Leiden
1:$G^{t-1},G^{t}$: Previous, current input graph
@@
8:$F$: Lambda functions passed to parallel Leiden (Alg. 4)
@@
10:function naiveDynamic($G^{t-1},G^{t},\Delta^{t-},\Delta^{t+},C^{t-1},K^{t-1},\Sigma^{t-1}$)
@@
1:$G^{t-1},G^{t}$: Previous, current input graph
@@
10:$F$: Lambda functions passed to parallel Leiden (Alg. 4)
@@
12:function deltaScreening($G^{t-1},G^{t},\Delta^{t-},\Delta^{t+},C^{t-1},K^{t-1},\Sigma^{t-1}$)
@@
A.3. Our Parallel Dynamic Frontier (DF) Leiden
Algorithm 3 presents the pseudocode for our Parallel Dynamic Frontier (DF) Leiden. It takes as input the previous graph snapshot $G^{t-1}$, the current graph snapshot $G^{t}$, edge deletions $\Delta^{t-}$ and insertions $\Delta^{t+}$ from the batch update, the previous community assignments $C^{t-1}$ for each vertex, the previous weighted degrees $K^{t-1}$ of vertices, and the previous total edge weights $\Sigma^{t-1}$ of communities. It outputs the updated community memberships $C^{t}$ for vertices, the updated weighted degrees $K^{t}$ of vertices, and the updated total edge weights $\Sigma^{t}$ of communities.
Algorithm 3 Our Parallel Dynamic Frontier (DF) Leiden (Sahu, 2024).
1:$G^{t-1},G^{t}$: Previous, current input graph
3:$C^{t-1},C^{t}$: Previous, current community of each vertex
4:$K^{t-1},K^{t}$: Previous, current weighted-degree of vertices
6:$\delta V$: Flag vector indicating if each vertex is affected
7:$isAffected(i)$: Is vertex $i$ is marked as affected?
8:$inAffectedRange(i)$: Can $i$ be incrementally marked?
9:$onChange(i)$: What happens if $i$ changes its community?
10:$F$: Lambda functions passed to parallel Leiden (Alg. 4)
11:
12:function dynamicFrontier($G^{t-1},G^{t},\Delta^{t-},\Delta^{t+},C^{t-1},K^{t-1},\Sigma^{t-1}$)
[…]
14:     for all $(i,j)\in\Delta^{t-}$ in parallel do
15:         if $C^{t-1}[i]=C^{t-1}[j]$ then $\delta V[i]\leftarrow 1$
16:     for all $(i,j,w)\in\Delta^{t+}$ in parallel do
17:         if $C^{t-1}[i]\neq C^{t-1}[j]$ then $\delta V[i]\leftarrow 1$
18:     function isAffected($i$)
19:         return $\delta V[i]$
20:     function inAffectedRange($i$)
21:         return $1$
22:     function onChange($i$)
23:         for all $j\in G^{t}.neighbors(i)$ do $\delta V[j]\leftarrow 1$
24:     $F\leftarrow\{isAffected,inAffectedRange,onChange\}$
[…]
28:     $C^{t}\leftarrow leiden(G^{t},C^{t-1},K^{t},\Sigma^{t},F)$
29:     return $\{C^{t},K^{t},\Sigma^{t}\}$
A.4. Our Dynamic-supporting Parallel Leiden
Algorithm 4 Our Dynamic-supporting Parallel Leiden (Sahu, 2023).
1:$G^{t}$: Current input graph
2:$C^{t-1}$: Previous community of each vertex
3:$K^{t}$: Current weighted-degree of each vertex
4:$\Sigma^{t}$: Current total edge weight of each community
5:$F$: Lambda functions passed to parallel Leiden
6:$G^{\prime}$: Current/super-vertex graph.
7:$C,C^{\prime}$: Current community of each vertex in $G^{t}$, $G^{\prime}$
8:$K,K^{\prime}$: Current weighted-degree of each vertex in $G^{t}$, $G^{\prime}$
9:Σ,$\Sigma,\Sigma^{\prime}$: Current total edge weight of each community in $G^{t}$, $G^{\prime}$
10:$\tau,\tau_{agg}$: Iteration, aggregation tolerance
11:
12:function leiden($G^{t},C^{t-1},K^{t},\Sigma^{t},F$)
[…]
14:     for all $i\in V^{t}$ do
15:         if $F.isAffected(i)$ then Mark $i$ as unprocessed
[…]
17:     Vertex membership: $C\leftarrow[0..|V^{t}|)$
18:     $G^{\prime}\leftarrow G^{t}$ ; $C^{\prime}\leftarrow C^{t-1}$ ; $K^{\prime}\leftarrow K^{t}$ ; $\Sigma^{\prime}\leftarrow\Sigma^{t}$
[…]
21:         $l_{i}\leftarrow leidenMove(G^{\prime},C^{\prime},K^{\prime},\Sigma^{\prime},F)$triangleright Alg. 5
22:         $C^{\prime}_{B}\leftarrow C^{\prime}$ ; $C^{\prime}\leftarrow[0..|V^{\prime}|)$ ; $\Sigma^{\prime}\leftarrow K^{\prime}$
23:         $l_{j}\leftarrow leidenRefine(G^{\prime},C^{\prime}_{B},C^{\prime},K^{\prime},\Sigma^{\prime},\tau)$triangleright Alg. 6
24:         if not first pass and $l_{i}+l_{j}\leq 1$ then break $\triangleright$ Done?
25:         $|\Gamma|,|\Gamma_{old}|\leftarrow$ Number of communities in $C$, $C^{\prime}$
26:         if $|\Gamma|/|\Gamma_{old}|>\tau_{agg}$ then break $\triangleright$ Low shrink?
27:         $C^{\prime}\leftarrow$ Renumber communities in $C^{\prime}$
28:         $C\leftarrow$ Lookup dendrogram using $C$ to $C^{\prime}$
29:         $G^{\prime}\leftarrow leidenAggregate(G^{\prime},C^{\prime})$triangleright Alg. 7
30:         $\Sigma^{\prime}\leftarrow K^{\prime}\leftarrow$ Weight of each vertex in $G^{\prime}$
31:         Mark all vertices in $G^{\prime}$ as unprocessed
32:         $C^{\prime}\leftarrow[0..|V^{\prime}|)$triangleright Use refine-based membership
[…]
34:     $C\leftarrow$ Lookup dendrogram using $C$ to $C^{\prime}$
35:     return $C$
In the algorithm, we begin by marking affected vertices as unprocessed (lines 14-15). Following this, the initialization phase initializes various parameters: the community membership of each vertex $C$ in $G^{t}$, the total edge weight of each vertex $K^{\prime}$, the total edge weight of each community $\Sigma^{\prime}$, and the community membership $C^{\prime}$ of each vertex in the current/super-vertex graph $G^{\prime}$ (lines 17-18). Subsequently, a series of passes (limited to $MAX\_PASSES$) comprising local-moving and aggregation phases are conducted (lines 20-33). Within each pass, the local-moving phase of the Leiden algorithm (Algorithm 5) is executed to optimize community assignments (line 21). Then, the community bound of each vertex is set for the refinement phase, and the memberships of vertices and the total weights of communities are reset as singleton vertices (line 22). The refinement phase, invoked by leidenRefine(), optimizes the community assignments of each vertex within its community bound (line 23). If the local-moving phase converges in a single iteration, indicating global convergence, the passes are terminated (line 24). Alternatively, if the drop in community count $|\Gamma|$ is deemed insignificant, indicating diminishing returns, the current pass is halted (line 26). If convergence conditions are not met, the aggregation phase is executed, involving community renumbering, updating top-level community memberships $C$ using dendrogram lookup, performing the aggregation phase (Algorithm 7), and adjusting the convergence threshold for subsequent passes via threshold scaling (line 33). The subsequent pass commences (line 20). After all passes, a final update of the top-level community membership $C$ of each vertex in $G^{t}$ is done via dendrogram lookup before returning it (lines 34-35).
A.4.1. Local-moving phase of our Parallel Leiden
The pseudocode for the local-moving phase is presented in Algorithm 5. It iteratively moves vertices among communities in order to maximize modularity. Here, the leidenMove() function operates on the current graph $G^{\prime}$, community membership $C^{\prime}$, total edge weight of each vertex $K^{\prime}$, total edge weight of each community $\Sigma^{\prime}$, and a set of lambda functions as inputs, yielding the number of iterations performed $l_{i}$.
Algorithm 5 Local-moving phase of our Parallel Leiden (Sahu, 2023).
1:$G^{\prime}$: Input/super-vertex graph
@@ A.4.1. Local-moving phase of our Parallel Leiden
5:$F$: Lambda functions passed to parallel Leiden
6:$H_{t}$: Collision-free per-thread hashtable
7:$l_{i}$: Number of iterations performed
8:$\tau$: Per iteration tolerance
9:
10:function leidenMove($G^{\prime},C^{\prime},K^{\prime},\Sigma^{\prime},F$)
@@
12:         Total delta-modularity per iteration: $\Delta Q\leftarrow 0$
13:         for all unprocessed $i\in V^{\prime}$ in parallel do
14:              Mark $i$ as processed (prune)
15:              if not $F.inAffectedRange(i)$ then continue
16:              $H_{t}\leftarrow scanCommunities(\{\},G^{\prime},C^{\prime},i,false)$
@@
18:              $c^{*}\leftarrow$ Best community linked to $i$ in $G^{\prime}$
19:              $\delta Q^{*}\leftarrow$ Delta-modularity of moving $i$ to $c^{*}$
20:              if $c^{*}=C^{\prime}[i]$ then continue
21:              $\Sigma^{\prime}[C^{\prime}[i]]-=K^{\prime}[i]$ ; $\Sigma^{\prime}[c^{*}]+=K^{\prime}[i]$ atomic
22:              $C^{\prime}[i]\leftarrow c^{*}$ ; $\Delta Q\leftarrow\Delta Q+\delta Q^{*}$
23:              Mark neighbors of $i$ as unprocessed
24:         if $\Delta Q\leq\tau$ then break $\triangleright$ Locally converged?
25:     return $l_{i}$
26:
27:function scanCommunities($H_{t},G^{\prime},C^{\prime},i,self$)
28:     for all $(j,w)\in G^{\prime}.edges(i)$ do
29:         if not $self$ and $i=j$ then continue
30:         $H_{t}[C^{\prime}[j]]\leftarrow H_{t}[C^{\prime}[j]]+w$
31:     return $H_{t}$
@@
Lines 11-24 encapsulate the primary loop of the local-moving phase. In line 12, we initialize the total delta-modularity per iteration $\Delta Q$. Subsequently, in lines 13-23, we concurrently iterate over unprocessed vertices. For each vertex $i$, we perform vertex pruning by marking $i$ as processed (line 14). Next, we verify if $i$ falls within the affected range (i.e., it is permitted to be incrementally marked as affected), and if not, we proceed to the next vertex (line 15). For each unskipped vertex $i$, we scan communities connected to $i$ (line 16), excluding itself, ascertain the optimal community $c*$ to move $i$ to (line 18), compute the delta-modularity of moving $i$ to $c*$ (line 19), update the community membership of $i$ (lines 21-22), and mark its neighbors as unprocessed (line 23) if a superior community is identified. It’s worth noting that this practice of marking neighbors of $i$ as unprocessed, which is part of the vertex pruning optimization, also aligns with algorithm of DF Leiden — which marks its neighbors as affected, when a vertex changes its community. Thus, vertex pruning facilitates incremental expansion of the set of affected vertices without requiring any extra code. In line 24, we examine whether the local-moving phase has achieved convergence (locally); if so, the loop is terminated (or if $MAX\_ITERATIONS$ is reached). Finally, in line 25, we return the number of iterations performed $l_{i}$.
A.4.2. Refinement phase of our Parallel Leiden
The pseudocode outlining the refinement phase of our Parallel Leiden is presented in Algorithm 5. This phase closely resembles the local-moving phase but incorporates the community membership obtained for each vertex as a community bound. In this phase, each vertex is required to select a community within its community bound to join, aiming to maximize modularity through iterative movements between communities, akin to the local-moving phase. At the onset of the refinement phase, the community membership of each vertex is reset so that each vertex initially forms its own community. The leidenRefine() function is employed, taking as input the current graph $G^{\prime}$, the community bound of each vertex $C^{\prime}_{B}$, the initial community membership $C^{\prime}$ of each vertex, the total edge weight of each vertex $K^{\prime}$, the initial total edge weight of each community $\Sigma^{\prime}$, and the current tolerance per iteration $\tau$, and returns the number of iterations executed $l_{j}$.
Algorithm 6 Refinement phase of our Parallel Leiden (Sahu, 2023).
1:$G^{\prime}$: Input/super-vertex graph
2:$C^{\prime}$: Community membership of each vertex
3:$K^{\prime}$: Total edge weight of each vertex
4:$\Sigma^{\prime}$: Total edge weight of each community
5:$G^{\prime}_{C^{\prime}}$: Community vertices (CSR)
6:$H_{t}$: Collision-free per-thread hashtable
7:$\tau$: Per iteration tolerance
8:
9:function leidenRefine($G^{\prime},C^{\prime}_{B},C^{\prime},K^{\prime},\Sigma^{\prime},\tau$)
10:     for all $i\in V^{\prime}$ in parallel do
11:         $c\leftarrow C^{\prime}[i]$
12:         if $\Sigma^{\prime}[c]\neq K^{\prime}[i]$ then continue
13:         $H_{t}\leftarrow scanBounded(\{\},G^{\prime},C^{\prime}_{B},C^{\prime},i,false)$
@@ A.4.2. Refinement phase of our Parallel Leiden
15:         $c^{*}\leftarrow$ Best community linked to $i$ in $G^{\prime}$ within $C^{\prime}_{B}$
16:         $\delta Q^{*}\leftarrow$ Delta-modularity of moving $i$ to $c^{*}$
17:         if $c^{*}=c$ then continue
18:         if $atomicCAS(\Sigma^{\prime}[c],K^{\prime}[i],0)=K^{\prime}[i]$ then
19:              $\Sigma^{\prime}[c^{*}]+=K^{\prime}[i]$ atomically
20:              $C^{\prime}[i]\leftarrow c^{*}$
21:
22:function scanBounded($H_{t},G^{\prime},C^{\prime}_{B},C^{\prime},i,self$)
23:     for all $(j,w)\in G^{\prime}.edges(i)$ do
24:         if not $self$ and $i=j$ then continue
25:         if $C^{\prime}_{B}[i]\neq C^{\prime}_{B}[j]$ then continue
26:         $H_{t}[C^{\prime}[j]]\leftarrow H_{t}[C^{\prime}[j]]+w$
27:     return $H_{t}$
28:
29:function atomicCAS($pointer,old,new$)
@@
31:     if $pointer=old$ then $pointer\leftarrow new$ ; return $old$
32:     else return $pointer$
Lines 10-20 embody the central aspect of the refinement phase. During this phase, we execute what is termed the constrained merge procedure (Traag et al., 2019). The essence of this procedure lies in enabling vertices, within each community boundary, to create sub-communities solely by permitting isolated vertices (i.e., vertices belonging to their own community) to alter their community affiliation. This process divides any internally-disconnected communities identified during the local-moving phase and prevents the emergence of new disconnected communities. Specifically, for every isolated vertex $i$ (line 12), we explore communities connected to $i$ within the same community boundary - excluding itself (line 13). Subsequently, we determine the optimal community $c*$ for relocating $i$ (line 15), and assess the delta-modularity of transferring $i$ to $c*$ (line 16). If a superior community is identified, we endeavor to update the community affiliation of $i$ provided it remains isolated (lines 18-20).
A.4.3. Aggregation phase of our Parallel Leiden

### v4

@@ A.1. Our Parallel Naive-dynamic (ND) Leiden
Algorithm 4 details our multicore implementation of Naive dynamic (ND) Leiden. Here, vertices are assigned to communities based on the previous snapshot of the graph, with all vertices being processed regardless of edge deletions and insertions in the batch update. The algorithm accepts the current/updated graph snapshot $G^{t}$, edge deletions $\Delta^{t-}$ and insertions $\Delta^{t+}$ in the batch update, the previous community membership $C^{t-1}$ for each vertex, the weighted degree of each vertex $K^{t-1}$, and the total edge weight of each community $\Sigma^{t-1}$. The output includes the updated community memberships $C^{t}$, the updated weighted degrees $K^{t}$, and the updated total edge weights of communities $\Sigma^{t}$.
@@
1:$G^{t}(V^{t},E^{t})$: Current/updated input graph
@@
8:$F$: Lambda functions passed to parallel Leiden (Alg. 2)
@@
10:function naiveDynamic($G^{t},\Delta^{t-},\Delta^{t+},C^{t-1},K^{t-1},\Sigma^{t-1}$)
@@
1:$G^{t}(V^{t},E^{t})$: Current/updated input graph
@@
10:$F$: Lambda functions passed to parallel Leiden (Alg. 2)
@@
12:function deltaScreening($G^{t},\Delta^{t-},\Delta^{t+},C^{t-1},K^{t-1},\Sigma^{t-1}$)
@@
A.2. Our Parallel Delta-screening (DS) Leiden
Algorithm 5 presents the pseudocode for our multicore implementation of Delta-screening (DS) Leiden. It employs modularity-based scoring to identify an approximate region of the graph where vertices are likely to change their community membership (Zarayeneh and Kalyanaraman, 2021). It takes as input the current/updated graph snapshot $G^{t}$, edge deletions $\Delta^{t-}$ and insertions $\Delta^{t+}$ in the batch update, the previous community memberships of vertices $C^{t-1}$, the weighted degrees of vertices $K^{t-1}$, and the total edge weights of communities $\Sigma^{t-1}$. The algorithm outputs the updated community memberships $C^{t}$, weighted degrees $K^{t}$, and total edge weights of communities $\Sigma^{t}$. Prior to processing, the batch update — which includes edge deletions $(i,j,w)\in\Delta^{t-}$ and insertions $(i,j,w)\in\Delta^{t+}$ — are sorted separately by the source vertex ID $i$.
@@ A.2. Our Parallel Delta-screening (DS) Leiden
A.3. Our Dynamic-supporting Parallel Leiden
A.3.1. Local-moving phase of our Parallel Leiden
The pseudocode for the local-moving phase is presented in Algorithm 6. It iteratively moves vertices among communities in order to maximize modularity. Here, the leidenMove() function operates on the current graph $G^{\prime}$, community membership $C^{\prime}$, total edge weight of each vertex $K^{\prime}$, total edge weight of each community $\Sigma^{\prime}$, changed communities flag vector $\Delta C^{\prime}$, and a set of lambda functions as inputs, yielding the number of iterations performed $l_{i}$.
Algorithm 6 Local-moving phase of our Parallel Leiden (Sahu et al., 2024a).
1:$G^{\prime}(V^{\prime},E^{\prime})$: Input/super-vertex graph
@@ A.3.1. Local-moving phase of our Parallel Leiden
5:$\Delta C^{\prime}$: Changed communities flag
6:$F$: Lambda functions passed to parallel Leiden
7:$H_{t}$: Collision-free per-thread hashtable
8:$l_{i}$: Number of iterations performed
9:$\tau$: Per iteration tolerance
10:
11:function leidenMove($G^{\prime},C^{\prime},K^{\prime},\Sigma^{\prime},\Delta C^{\prime},F$)
@@
13:         Total delta-modularity per iteration: $\Delta Q\leftarrow 0$
14:         for all unprocessed $i\in V^{\prime}$ in parallel do
15:              Mark $i$ as processed (prune)
16:              if not $F.inAffectedRange(i)$ then continue
17:              $H_{t}\leftarrow scanCommunities(\{\},G^{\prime},C^{\prime},i,false)$
@@
19:              $c^{*}\leftarrow$ Best community linked to $i$ in $G^{\prime}$
20:              $\delta Q^{*}\leftarrow$ Delta-modularity of moving $i$ to $c^{*}$
21:              if $c^{*}=C^{\prime}[i]$ then continue
22:              $\Sigma^{\prime}[C^{\prime}[i]]-=K^{\prime}[i]$ ; $\Sigma^{\prime}[c^{*}]+=K^{\prime}[i]$ atomic
23:              $C^{\prime}[i]\leftarrow c^{*}$ ; $\Delta Q\leftarrow\Delta Q+\delta Q^{*}$
24:              Mark neighbors of $i$ as unprocessed
25:              if is dynamic alg. then $\Delta C^{\prime}[c]\leftarrow\Delta C^{\prime}[c^{*}]\leftarrow 1$
26:         if $\Delta Q\leq\tau$ then break $\triangleright$ Locally converged?
27:     return $l_{i}$
28:
29:function scanCommunities($H_{t},G^{\prime},C^{\prime},i,self$)
30:     for all $(j,w)\in G^{\prime}.edges(i)$ do
31:         if not $self$ and $i=j$ then continue
32:         $H_{t}[C^{\prime}[j]]\leftarrow H_{t}[C^{\prime}[j]]+w$
33:     return $H_{t}$
@@
Lines 12-26 encapsulate the primary loop of the local-moving phase. In line 13, we initialize the total delta-modularity per iteration $\Delta Q$. Subsequently, in lines 14-25, we concurrently iterate over unprocessed vertices. For each vertex $i$, we perform vertex pruning by marking $i$ as processed (line 15). Next, we verify if $i$ falls within the affected range (i.e., it is permitted to be incrementally marked as affected), and if not, we proceed to the next vertex (line 16). For each unskipped vertex $i$, we scan communities connected to $i$ (line 17), excluding itself, ascertain the optimal community $c*$ to move $i$ to (line 19), compute the delta-modularity of moving $i$ to $c*$ (line 20), update the community membership of $i$ (lines 22-23), and mark its neighbors as unprocessed (line 24) if a superior community is identified. It is worth noting that this practice of marking neighbors of $i$ as unprocessed, which is part of the vertex pruning optimization, also aligns with algorithm of DF Leiden — which marks its neighbors as affected, when a vertex changes its community. Thus, vertex pruning facilitates incremental expansion of the set of affected vertices without requiring any extra code. Further, for ND, DS, and DF Leiden, we mark the source $c$ and target communities $c^{*}$ of the migrating vertex $i$ as changed. With Static Leiden, all communities are always refined. In line 26, we examine whether the local-moving phase has achieved convergence (locally); if so, the loop is terminated (or if $MAX\_ITERATIONS$ is reached). Finally, in line 27, we return the number of iterations performed $l_{i}$.
A.3.2. Refinement phase of our Parallel Leiden
Algorithm 7 Refinement phase of our Parallel Leiden (Sahu et al., 2024a).
1:$G^{\prime}(V^{\prime},E^{\prime})$: Input/super-vertex graph
2:$C^{\prime}_{B}$: Community bound of each vertex
3:$C^{\prime}$: Community membership of each vertex
4:$K^{\prime}$: Total edge weight of each vertex
5:$\Sigma^{\prime}$: Total edge weight of each community
6:$\Delta C^{\prime}$: Community changed flag
7:$G^{\prime}_{C^{\prime}}$: Community vertices (CSR)
8:$H_{t}$: Collision-free per-thread hashtable
9:$\tau$: Per iteration tolerance
10:
11:function leidenRefine($G^{\prime},C^{\prime}_{B},C^{\prime},K^{\prime},\Sigma^{\prime},\Delta C^{\prime},\tau$)
12:     for all $i\in V^{\prime}$ in parallel do
13:         $c\leftarrow C^{\prime}[i]$
14:         if $\Delta C^{\prime}[c]=0$ or $\Sigma^{\prime}[c]\neq K^{\prime}[i]$ then continue
15:         $H_{t}\leftarrow scanBounded(\{\},G^{\prime},C^{\prime}_{B},C^{\prime},i,false)$
@@ A.3.2. Refinement phase of our Parallel Leiden
17:         $c^{*}\leftarrow$ Best community linked to $i$ in $G^{\prime}$ within $C^{\prime}_{B}$
18:         $\delta Q^{*}\leftarrow$ Delta-modularity of moving $i$ to $c^{*}$
19:         if $c^{*}=c$ or $C^{\prime}[c^{*}]\neq c^{*}$ then continue
20:         if $atomicCAS(\Sigma^{\prime}[c],K^{\prime}[i],0)=K^{\prime}[i]$ then
21:              $\Sigma^{\prime}[c^{*}]+=K^{\prime}[i]$ atomically
22:              $C^{\prime}[i]\leftarrow c^{*}$
23:
24:function scanBounded($H_{t},G^{\prime},C^{\prime}_{B},C^{\prime},i,self$)
25:     for all $(j,w)\in G^{\prime}.edges(i)$ do
26:         if not $self$ and $i=j$ then continue
27:         if $C^{\prime}_{B}[i]\neq C^{\prime}_{B}[j]$ then continue
28:         $H_{t}[C^{\prime}[j]]\leftarrow H_{t}[C^{\prime}[j]]+w$
29:     return $H_{t}$
30:
31:function atomicCAS($pointer,old,new$)
@@
33:     if $pointer=old$ then $pointer\leftarrow new$ ; return $old$
34:     else return $pointer$
The pseudocode outlining the refinement phase of our Parallel Leiden is presented in Algorithm 6. This phase closely resembles the local-moving phase but incorporates the community membership obtained for each vertex as a community bound. In this phase, each vertex is required to select a community within its community bound to join, aiming to maximize modularity through iterative movements between communities, akin to the local-moving phase. At the onset of the refinement phase, the community membership of each vertex is reset so that each vertex initially forms its own community. The leidenRefine() function is employed, taking as input the current graph $G^{\prime}$, the community bound of each vertex $C^{\prime}_{B}$, the initial community membership $C^{\prime}$ of each vertex, the total edge weight of each vertex $K^{\prime}$, the initial total edge weight of each community $\Sigma^{\prime}$, changed communities flag vector $\Delta C^{\prime}$, and the current tolerance per iteration $\tau$, and returns the number of iterations executed $l_{j}$.
Lines 12-22 embody the central aspect of the refinement phase. During this phase, we execute what is termed the constrained merge procedure (Traag et al., 2019). The essence of this procedure lies in enabling vertices, within each community boundary, to create sub-communities solely by permitting isolated vertices (i.e., vertices belonging to their own community) to alter their community affiliation. This process divides any internally-disconnected communities identified during the local-moving phase and prevents the emergence of new disconnected communities. Specifically, for every isolated vertex $i$ (line 14), we explore communities connected to $i$ within the same community boundary - excluding itself (line 15). The refinement phase is skipped for a vertex if its community has not been flagged as changed ($\Delta C^{\prime}[c]=0$), or if another vertex has joined its community, as indicated by the total community weight no longer matching the total edge weight of the vertex. Subsequently, we determine the optimal community $c*$ for relocating $i$ (line 17), and assess the delta-modularity of transferring $i$ to $c*$ (line 18). If a superior community is identified, we attempt to update the community affiliation of $i$ provided it remains isolated (lines 20-22). Note that we do not migrate the vertex $i$ to community $c^{*}$ if the vertex representing $c^{*}$ has itself moved to another community.
A.3.3. Aggregation phase of our Parallel Leiden